What will be the approximate population of the United States, if its current population of 316,000,000 grows at a compound rate of 2.3% annually for 25 years? (Enter your answer rounded to the nearest whole number; for example, record 218,934,210.86 as 218934211).
Answer #1

This question requires application of basic time value of money function, according to which FV = PV * (1 + r) where FV is F

FV = 316,000,000 * (1 + 2.3%)25

FV = 316,000,000 * 1.76559399

FV = 557,927,702